# AgentHub: A Research Agenda for Agent Sharing Infrastructure

### v1

@@ 3.2.2. Lifecycle Transparency
Future work should define lifecycle metadata standards with clear states, timestamps, and rationales, along with monitoring to detect abandonment or unexpected behavioral drift. Research must also address how deprecated or revoked agents should appear in discovery, who has authority to mark or revoke them, and how federated mirrors should coordinate state changes. Comparing agent lifecycles with traditional software lifecycles may reveal where new loops–such as agents participating in design and implementation–demand stronger provenance, transparency, and control.
3.2.3. Ecosystem Interoperability
One Schema, Many Dialects: Descriptors in MCP (tools), A2A (roles/behaviors), and ACP (message schemas/policies) use different primitives and evolve at different speeds. Without a standardized, signed manifest, semantics may be lost in translation, caches drift across mirrors, and naive popularity metrics dominate cross-protocol rankings while ignoring evidence quality and freshness. Stable identifiers for referenced artifacts are also missing, making cross-registry links fragile.
Short-term priorities include a compact capability ontology and a canonical manifest with required fields (capabilities, I/O modalities, protocol bindings, permissions, SBOM-style dependencies) and optional per-ecosystem extensions. Declarative adapters can then map native descriptors to this core and back, validated via round-trip conformance tests. Stable cross-registry identifiers (e.g., Purl-style) should be introduced and tested end-to-end with npm, PyPI, and model hubs. Early benchmarks could measure cross-protocol discovery in terms of precision/recall for intent queries, preservation of evidence link, freshness under churn, and ranking fairness as ecosystems scale.
3.2.4. Openness and Governance
Workforce and Concentration of Expertise: Many projects are maintained by a single person, creating bus-factor risk. Zimmermann et al. emphasize that the issue is less maintainer shortage than concentrated control (Zimmermann2019, ). Governance models differ: npm allows instant publication, while curated ecosystems like CRAN impose stricter checks.
Future work should investigate governance structures that combine openness with verifiability. Possible directions include decentralized namespace assignment rooted in ANS, community-driven policy boards for resolving disputes, and auditable logs of publication and revocation decisions. Comparative studies of centralized versus federated governance models can illuminate trade-offs in consistency, adoption, and resilience. Open questions include how to embed transparency in decision-making without sacrificing agility, and how to make governance mechanisms both fair across domains and enforceable at internet scale.
3.2.5. Trust and Security
Privacy and Vulnerability: In software registries, account hijacks and malicious updates already erode trust (Zimmermann2019, ). LLM-based agents add privacy-specific attack surfaces: misconfiguration or weak governance can leak sensitive data or intellectual property (wang-etal-2025-unveiling-privacy, ; zharmagambetov2025agentdamprivacyleakageevaluation, ); training-data exposure can reconstruct confidential content (carlini2021extractingtrainingdatalarge, ); and real incidents have leaked corporate data (samsung_chatgpt_forbes_2023, ). Multi-agent prompt injection can propagate malicious instructions and compromise collective decision-making (lee2024promptinfectionllmtollmprompt, ; shi2025promptinjectionattacktool, ). Privilege escalation, guardrail bypasses, weak output validation, and insufficient access control enable unauthorized tool use and code execution (narajala2025securinggenaimultiagentsystems, ). The attack frontier remains underexplored.
AgentHub can deploy defenses such as signed manifests, provenance attestations (pypi-pep458, ; npm-trusted, ), verified namespaces via ANS (ans, ), and protections against typosquatting and package-confusion attacks (Zimmermann2019, ; jiang2025confuguardusingmetadatadetect, ). AI-specific risks include deserialization exploits and prompt-injection. It is crucial to test whether defenses from human-mediated registries hold for autonomous agent systems. New threats call for zero-trust privilege separation, runtime checks for I/O behavior, and privacy-preserving audit pipelines to verify evidence without leaks (he2025redteamingllmmultiagentsystems, ; narajala2025securinggenaimultiagentsystems, ; samsung_chatgpt_forbes_2023, ; wang-etal-2025-unveiling-privacy, ).
3.2.6. Discovery and Workflow Integration
Keyword Search and Popularity Bias: Agents that recommend or install one another can create feedback loops that instill mediocre or unsafe entries. Experience from software registries highlights how naive signals mislead: Zerouali et al. show that different popularity metrics in npm yield inconsistent results (Zerouali2019, ), underlining how reliance on downloads or stars can mislead users; and Jones et al. observed that model popularity on HuggingFace correlates strongly with documentation quality (jones2024knowhuggingfacesystematic, ), suggesting that discovery in AgentHub would similarly benefit from incentives for high-quality documentation. Humans can partly correct for this by reading docs. Agents benefit greatly if those signals are captured as structured metadata and verifiable evidence. Without workflow integration, both humans and agents may fall back to ad-hoc search, undermining adoption and reproducibility.
AgentHub can improve current practice by combining keyword search with structured metadata and evidence linked to benchmarks, incorporating lifecycle states to avoid unsafe or outdated agents. Evaluation should identify which signals such as metadata and documentation quality ensure reliable discovery. Discovery should match user goals to agents by capabilities and evidence, ensuring that popularity does not overshadow quality. This resembles work on recommendation systems for software libraries like Code Librarian, which uses contextual signals to suggest packages (tao2023codelibrarian, ). Research benchmarking discovery is necessary to measure precision, recall, and resilience under ecosystem changes.
4. Future Plans
4.1. Roadmap and Evaluation
We suggest that the community prioritize first the design of AgentHub’s core mechanisms: standardized capability schemas and evidence pipelines, lifecycle metadata for transparency, canonical manifests for interoperability, and basic governance structures. We should seed AgentHub with a small, diverse set of agents and provide a lightweight CLI/SDK so discovery fits ordinary workflows. Evaluation can emphasize usability and reliability, while also exploring which metrics are most appropriate. Security efforts can focus first on known risks such as typosquatting, prompt injection, and account compromise.
Longer term, the community must tackle challenges such as detecting and mitigating emergent multi-agent attacks, ensuring dynamic provenance and lifecycle accountability, and integrating AgentHub with broader software and model registries. Governance may need to evolve toward federated models with auditable processes, while discovery might mature into recommendation systems that balance intent, evidence quality, and fairness.
To support continuous measurement and adaptation, the operators of AgentHub should provide infrastructure to collect user feedback and registry-level longitudinal data, enabling future studies of common problems and ecosystem evolution. As AI supply chain standardization advances (bennet2025SPDXAIBOM, ; owasp-aibom, ), AgentHub should remain aligned with emerging specifications such as AIBOM. In parallel, as the agentic ecosystem expands, the AgentHub operators will refine the interface and reinforce integration and access controls to ensure secure, seamless interaction with new agents.
4.2. Positioning AgentHub if AGI Centralizes Orchestration
Our analysis of AgentHub is grounded in the current landscape and capabilities of agents. No one knows the future, but we acknowledge that increasingly capable foundation models may internalize a great deal of the coordination and planning that motivates today’s agentic and multi-agent systems.
We view this not as an end to multi-agent infrastructure but as a shift in where coordination lives and, importantly, a shift that heightens the need for transparent capability schemas, lifecycle visibility, and governance (section 3.2). First, economics and operational constraints make a purely “AGI-first” strategy brittle. Highly general models are likely to remain expensive in compute, energy, and latency and they will not be uniformly deployable across every environment that needs automation. Many production workflows will continue to mix a capable planner with specialized, composable executors that are cheaper to run, easier to colocate with data, and better aligned with SLA and locality requirements. If AgentHub exposes cost, latency, and fidelity signals alongside capabilities and evidence, discovery can optimize for fitness under budget and policy, rather than only for raw model capacity.

### v2

@@ 3.2.5. Lifecycle Transparency
Solution: Future work should define lifecycle metadata standards with clear states, timestamps, and rationales, along with monitoring to detect abandonment or unexpected behavioral drift. Research must also address how deprecated or revoked agents should appear in discovery, who has authority to mark or revoke them, and how federated mirrors should coordinate state changes. Comparing agent lifecycles with traditional software lifecycles may reveal where new loops–such as agents participating in design and implementation–demand stronger provenance, transparency, and control.
3.2.6. Capability Clarity & Evidence
Challenge: Fragmentation & Duplication. As npm’s “micro-packages” created brittle dependency chains (kula2017impactmicropackagesempiricalstudy, ; Abdalkareem2017TrivialPackages, ), agent ecosystems may spawn “micro-agents” with overlapping functions. Because agents can autonomously query registries and compose others, duplication and fragility can spread dynamically, enlarging the attack surface and degrading reliability, especially at scale as agents evolve and reimplement overlapping capabilities.
Challenge: Equivalence. Beyond functional duplication, a deeper challenge is determining when two agents are truly equivalent. At a syntactic level, the same agent may appear across multiple repositories, creating confusion about which copy is authoritative. At a semantic level, multiple agents may claim the same capabilities but diverge in behavior due to nondeterminism and evolution in models, changes in context, and even hardware selection. Addressing this requires more than metadata alignment: AgentHub should support persistent identifiers, cross-registry attestations, and re-executable evidence pipelines to assess whether two agents are truly equivalent.
Challenge: Abuse of Ambiguity. Ambiguity and missing metadata enable attacks in software registries (JiangPTMReuse, ); agents face the same risk. Unclear manifests let adversaries mimic popular entries or collaborators, echoing typosquatting and account hijackings in npm (Zimmermann2019, ). Agents’ autonomous installation/generation can accelerate such propagation unless strong signing and evidence requirements are enforced.
Solution: AgentHub will (i) standardize machine-readable capability schemas covering capabilities, modalities, protocol roles, and dependencies enforced at publish time (JiangPTMReuse, ; Jiang2025Naming, ); (ii) support lightweight, re-executable (idempotent) evidence pipelines linking claims to traces or benchmark runs across versions; and (iii) add badges or similar checks as digital nudges (brown2019nudges, ).
4. Prototype
AgentHub is grounded in a concrete, open-source reference implementation that makes the agenda’s mechanisms directly testable. The prototype is intentionally not an agent runtime. Instead, it is a registry layer that supports publishing, discovery, and reuse of agents under machine-checkable contracts, such as standardized capability manifests and evidence, lifecycle transparency, protocol interoperability, and a baseline trust and security model suitable for automated reuse. The goal of the reference implementation is to provide a minimal but complete substrate that can publish, index, search, resolve, verify, and evolve so that the requirements in Figure 2 can be evaluated empirically and iterated with the community.
4.1. Architecture & Scope
The prototype follows a design that stresses mutable metadata and immutable artifacts. Agent entries are represented as versioned snapshots of a canonical manifest stored immutably by content hash, while the registry maintains mutable metadata (indexes, lifecycle state, and evidence summaries) that can evolve without changing the underlying artifact. This makes verification straightforward, so clients can recompute the hash of a retrieved manifest and compare it to the registry’s recorded digest, and signatures can be defined over that digest rather than over an ambiguous, mutable document.
At a high level, the system exposes a small set of operations that map directly to the user and agent workflows described in section 3.2.1: (i) publish an agent version, (ii) search and rank candidates, (iii) resolve a stable identifier and version range to a concrete version and endpoint bindings, (iv) verify provenance signals, (v) attach evidence records, and (vi) append lifecycle events that affect discovery and resolution. Discovery is treated as a first-class interface, because in an agent ecosystem the users of the registry include both humans and agents, and discovery must integrate into planning and orchestration loops rather than requiring manual browsing.
4.2. Canonical Manifest, Evidence Records, and Lifecycle Events
A central deliverable of the prototype is a core manifest with publish-time validation. The manifest provides the minimum required fields that enable automation and reproducibility across heterogeneous ecosystems, including stable identifiers; declared capabilities and I/O modalities; protocol bindings; runtime permissions and preconditions; SBOM-style dependency references; explicit lifecycle state; and an evidence policy that declares what kinds of evidence are expected and how freshness is interpreted. Optional per-ecosystem extensions allow native descriptors to be carried without forcing premature convergence. This design follows our ”one schema, many dialects” requirement, as we will keep a required core that is stable enough to be validated and indexed, while preserving protocol-specific descriptors in an extension space (Figure 3).
Figure 3. Zoomed excerpt of an example AgentHub manifest illustrating capability declarations (I/O schemas and constraints) and protocol bindings for multiple ecosystems (A2A, MCP). AgentHub validates and indexes these fields to support intent-accurate discovery and cross-protocol resolution.
At publish time, the core manifest is validated under a strict schema, canonicalized deterministically, and stored as an immutable blob keyed by its hash. The schema is designed to be machine-actionable with capabilities including stable identifiers and, where possible, structured input/output expectations; protocol bindings name the interaction dialect (e.g., MCP vs. A2A) and provide resolvable endpoints; permissions declare resource access explicitly (network, filesystem, tool access, data handling), so that automated selection can incorporate risk constraints rather than relying on prose.
The prototype also defines an evidence record as a separate, version-bound object that links a capability claim (or a broader ”fitness claim”) to auditable artifacts. Each record binds to a specific agent version and manifest digest, and carries (i) the method/recipe used to produce the evidence (e.g., harness version, container digest/workflow reference), (ii) inputs (config, seeds, environment), (iii) outputs as immutable artifacts with hashes, and (iv) optional attestations. Evidence is stored immutably and referenced from the registry, enabling third parties to reproduce or independently re-run checks. This operationalizes our argument that agents benefit when discovery signals are captured as structured metadata and verifiable evidence rather than popularity or documentation alone.
Finally, the prototype treats lifecycle as an append-only event log rather than a single mutable flag. Agent versions move through explicit states (e.g., active, deprecated, retired, revoked), and each state transition is recorded with a timestamp and rationale. The registry’s default behaviors respect these states. Resolution must not return revoked/retired versions, and discovery must incorporate lifecycle state into ranking and filtering. This is essential for lifecycle visibility and fast revocation in ecosystems where reuse is automated and version churn is frequent.
4.3. Interoperability and Discovery as Registry Mechanisms
Interoperability in our prototype will be implemented as declaractive adapters that map native protocol descriptors into the core manifest and back. MCP and A2A use different primitives and evolve at different speeds, so a standardized, signed manifest is needed to prevent semantic loss and cache drift across mirrors. In the prototype, protocol bindings provide a stable, comparable surface (capabilities, modalities, endpoints, auth requirements), while the original descriptor (e.g., an A2A Agent Card or MCP tool descriptor) can be stored in extensions for lossless round-trip translation where possible. The concrete outcome is testable: adapters are validated through round-trip conformance tests that check preservation of core semantics and the stability of cross-registry identifiers.
Discovery is implemented as a registry concern because it is one of the primary ways automation will consume AgentHub. Our prototype will provide programmatic search over manifest fields and evidence signals, and a resolution interface that maps stable identifiers and version ranges to an exact version and protocol endpoints. Ranking is designed to be evidence forward. Rather than defaulting to popularity proxies, the registry returns candidates with explicit reasons such as relevance from structured metadata, evidence coverage and freshness, lifecycle state, and compatibility constraints. This directly responds to the ”keyword search and popularity bias” challenge identified in Figure 2.
5. Preliminary Evaluation
AgentHub’s research agenda (Figure 2) frames six requirements that collectively target four objectives: fair, intent-accurate discovery; resilient federated operations; cross-ecosystem portability; and reproducible, auditable reuse. In addition to early evaluations on discovery (discussed below), we propose a concrete evaluation program that combines (i) controlled benchmarks that isolate mechanisms and quantify trade-offs, (ii) adversarial experiments that test whether defenses remain effective under autonomous reuse, and (iii) longitudinal, deployment-level measurements that capture governance and ecosystem dynamics. These evaluation plans are discussed below.
5.1. Evaluating Discovery and Workflow Integration
5.1.1. Method:
We implemented a two-stage recommendation pipeline on top of ANS that recommends A2A agents using their Agent Cards as structured, comparable descriptions. Figure 4 summarizes the interaction pattern, where a client submits a query, the recommendation system consults the registry to retrieve and rank candidates, and the selected agent is returned to the client.
Figure 4. Two-stage discovery workflow used in our initial evaluation. A client query is handled by a recommendation system that consults the registry for candidate agents, then returns a ranked selection to the client.
In stage one, the recommendation system applies text-based retrieval over A2A agent cards to identify candidate agents relevant to a user query. We evaluate three retrieval strategies: a lexical approach (BM25), a semantic embedding-based approach, and a hybrid method that combines lexical and semantic signals. (8442792, ; bm25s, ; sultania2024domainspecificquestionansweringhybrid, ) This stage is designed to efficiently narrow the candidate set using static, declarative agent descriptions.
In stage two, the highest-ranked candidates from stage one are further evaluated using a structured, interview-based verification process. This phase engages agents directly to elicit behavioral evidence, allowing the system to assess whether an agent’s actual capabilities align with the inferred intent of the query. Importantly, this stage is intended to produce explicit evidence of suitability, rather than re-rank candidates based on textual similarity.
The evaluation registers eight A2A-based agents as live servers, meaning each agent is deployed as an independently running service that exposes its A2A interface and responds to real-time queries. The agents are intentionally organized into four pairs with overlapping but distinct capabilities to create controlled ambiguity during discovery (e.g., multiple debugging or performance-related agents with different specializations). We evaluate the system using 24 semi-ambiguous natural-language queries, each designed so that more than one agent could plausibly address the request, but only one agent represents the intended primary capability. For example, the query “Agent that diagnoses backend crashes occurring only under concurrent load” could be handled by several debugging-related agents, but the “code-debugging-assistant” is designated as the best match because it explicitly targets concurrency-related logic faults. Ground-truth labels for all queries were assigned based on domain expertise and each agent’s intended functional scope. This dataset design reflects realistic discovery scenarios in which advertised capabilities overlap, requiring deeper behavioral validation beyond surface-level descriptions.
The study measures ranking quality and efficiency using Precision@1, Recall@3, and latency, averaged across multiple runs to account for LLM nondeterminism in the interview stage. Precision@1 measures the fraction of queries for which the top-ranked (recommended) agent is the correct one. Recall@3 measures the fraction of queries for which the correct agent appears within the top three results. Latency captures the total end-to-end time required to process a query and return a recommendation. We first compare retrieval-only performance across lexical, semantic, and hybrid methods, and then evaluate the full two-phase pipeline that supplements retrieval with structured interview-based verification.
5.1.2. Results:
Hybrid retrieval achieves the best balance among retrieval-only methods (Precision@1 83.33%, Recall@3 95.83%, latency 0.57s), while the full pipeline maintains high Recall@3 (98.61% $\pm$ 1.97%) at substantially higher latency (101.30s $\pm$ 7.25s), explicitly positioning the interview stage as an evidence-producing mechanism rather than pure ranking booster. Table 2 summarizes the results of different phases and retrieval methods.
Table 2. Phase 1 Retrieval vs. Two-Phase Pipeline
Method P@1 R@3 Lat (s)
Phase 1: Text Retrieval
BM25 75.00 91.67 0.01
Semantic 66.67 100.00 0.85
Hybrid 83.33 95.83 0.57
Phase 2: Hybrid + Interviews
Full Pipeline 73.61 ± 1.97 98.61 ± 1.97 101.30 ± 7.25
6. Next Steps and Roadmap
We treat the results from section 5 as an initial benchmark for the ”Discovery & Workflow Integration” requirement, and we will expand it in three ways. First, we will scale beyond a small, curated pool into a seed corpus that includes heterogeneous agent types (tool-using agents, agent-to-agent coordinators, domain-specialized micro agents) and deliberately introduce documentation and metadata variance, reflecting real registry messiness rather than uniform A2A-only artifacts. Second, we will evaluate discovery not only by ranking metrics but also by downstream task success in workflows (e.g., whether a selected agent completes a CI/CD task under policy constraints), measuring end-to-end correctness and the cost of mis-selection. Third, we will evaluate robustness and fairness by injecting popularity signals and churn, testing whether intent and evidence can dominate naive ”stars/downloads” effects as ecosystems scale, consistent with the paper’s warning that popularity loops can dominate discovery.
6.1. Evaluating Trust and Security
AgentHub argues that autonomous composition widens the attack surface and makes provenance, revocation, and enforceable identity central. The key evaluation question is not whether individual defenses exist, but whether they hold when (a) agents can publish agents, (b) agents can select agents, and (c) reuse is automated at scale. We therefore will evaluate Trust & Security through an explicit registry red-team suite that attempts to publish and propagate adversarial artifacts, including typosquatting and package-confusion analogs, identity spoofing, over-privileged manifests, malicious updates, and prompt-injection or tool-misuse behaviors embedded in agent workflows.
Evaluation outcomes will include attack success rates under different enforcement policies (e.g., publish-time schema enforcement, signature requirements, namespace verification, evidence-gated promotion), time-to-detection and time-to-containment, and the degree to which security signals can be made machine-actionable without turning governance into opaque central control. We will also explicitly test for agents that misrepresent capabilities for selection advantage, as this was an observed weakness during our discovery evaluation. We will introduce adversarial capability inflation agents and measure whether evidence requirements and behavioral checks can detect and penalize strategic misrepresentation.
6.2. Evaluating Openness and Governance
The paper motivates governance as a balance between openness (low-friction publishing, transparent namespaces) with safeguards against spam, squatting, and opaque takedowns. We will therefore structure evaluation for this requirement in two phases. In the prototype phase, we run controlled publishing studies that measure submission friction, correctness of manifests under enforcement, and how often policy checks block legitimate submissions versus adversarial or low-quality ones. In the deployment phase, we will instrument the registry for longitudinal metrics such as submission volume, acceptance/rejection rates by reason code, median time-to-resolution for disputes, rates of automated publishing versus human publishing, and the concentration of publishing authority across namespaces.
We will also run governance stress tests that simulate realistic failure modes such as coordinated spam attempts, namespace disputes, and emergency revocations. The goal is to empirically characterize the trade-offs between curated and open models for agents, under the distinct condition that many publishers may themselves be automated pipelines.
6.3. Evaluating Ecosystem Interoperability
In AgentHub, we see interoperability as more than just supporting MCP and A2A; it is specifically ”one schema, many dialects,” with canonical manifests and declarative adapters that will be validated by round-trip conformance tests so that semantics are preserved across protocol translations. This requirement is well-suited to a rigorous, testable evaluation in which we will define a suite of reference agents expressed natively in different ecosystems (e.g., MCP tool descriptors and A2A agent cards), translate them into the AgentHub core manifest, and translate back, measuring round-trip fidelity at the level of capabilities, modalities, permissions, dependencies, and evidence links. We will complement this with cross-protocol discovery benchmarks, such as identical intent queries that should retrieve comparable agents regardless of whether the underlying descriptor dialect is MCP- or A2A-native. Evidence links should also remain resolvable and correctly bound to versions across protocol boundaries.
Interoperability evaluation will also include failure characterization to investigate which fields are lossy under translation, which protocol features cannot be expressed in a shared core without distortion, and what minimal portable contract is sufficient for safe cross-ecosystem reuse.
6.4. Evaluating Lifecycle Transparency
Lifecycle transparency becomes meaningful only if lifecycle state changes propagate quickly enough to affect automated reuse. This paper argues for explicit states (active, deprecated, rotated, retired, revoked) with timestamps and rationales, and for coordination across federated mirrors. We will evaluate this requirement via churn experiments that simulate realistic ecosystem dynamics including frequent version releases, dependency updates, emergency revocations, and key rotations. We will measure propagation latency of lifecycle events across mirrors and caches, the degree to which clients honor lifecycle state in discovery and selection, and failure cases where stale mirrors lead to unsafe reuse.
6.5. Evaluating Capability Clarity and Evidence
Capability clarity and evidence is the requirement that most sharply differentiates AgentHub from existing directories. Our near-term priorities include enforcing machine-readable capability schemas at publish time and supporting lightweight, re-executable evidence pipelines that link capability claims to benchmark runs or traces across versions. We will evaluate this along two axes: whether manifests become sufficiently precise to support automated planning and risk checks, and whether evidence is both (a) discriminative for selection and (b) reproducible and auditable.
On precision, we plan to test whether an enforced manifest schema reduces ambiguity and capability inflation by measuring inter-rater agreement among humans and agents when mapping natural-language intents to manifest claims, and by measuring how often over-privileged or under-specified agents can pass publish-time validations. On evidence, we will treat evidence records as structured, signed, re-runnable artifacts and quantify (i) reproducibility across environments and repeated runs, (ii) stability under model or dependency updates, and (iii) the relationship between evidence signals and downstream task success. Interview-based verification can be used as one evidence mechanism for ambiguous intents, and we will assess its value in terms of confidence/correctness trade-offs.
6.6. Practical Considerations
Beyond ranking quality, we will report operational costs that determine deployability. We will measure (i) publish-time overhead for validation, canonicalization, and signature checks, (ii) end-to-end query latency for search, resolution, and verification, and (iii) storage growth for manifests, evidence artifacts, and lifecycle logs. We will scale the number of registered agents and update churn to report throughput (publishes, queries), tail latency, and the dominant bottlenecks.
6.7. Roadmap (Jan–Sep 2026).
We treat this workshop paper as Phase 1 of a longer effort that culminates in a journal-length contribution and a reusable artifact suite. Over the next eight months we will stabilize the core contract, strengthen interoperability and evidence plumbing, expand evaluation coverage across all six requirements, and release artifacts that others can adopt and extend. Jan–Mar: freeze the v0.x manifest and evidence schemas; implement publish-time validation and canonicalization; add lifecycle events and signature verification; release a minimal CLI/API. Apr–May: implement declarative adapters for at least two dialects (MCP and A2A) with round-trip conformance tests; evaluate lifecycle freshness and revocation propagation under churn and caching. Jun–Aug: scale the seed corpus and intent suite; broaden from discovery metrics to workflow-level outcomes under policy constraints; add a robustness suite (inflation, malicious updates, typosquatting, prompt/tool misuse). Sep: package schemas, validators, adapters, benchmarks, and harnesses as a reproducible bundle and submit the journal-length manuscript.
7. CONCLUSION
Agent ecosystems are scaling faster than the infrastructure needed to make them discoverable, comparable, and governable under automated reuse. We advance AgentHub as a registry substrate and research agenda grounded in a reference implementation centered on (i) a canonical, publish-time validated manifest that is indexable yet extensible, (ii) version-bound evidence records that connect capability and fitness claims to auditable artifacts, and (iii) lifecycle states enforced by default in resolution and discovery. Our discovery benchmark quantifies the value of structured, machine-actionable metadata for retrieval and clarifies when behavioral evidence is needed under ambiguous intents, motivating the broader evaluation program in section 5. Together, AgentHub moves selection and governance signals out of prose and into enforceable, machine-checkable contracts, enabling auditable, comparable, and automatable agent reuse.
Data Availability. The artifact, including manifest.yaml and evidence_record.yaml, is available at (agenthub-artifact, ).